### v1

# Efficient Batch and Recursive Least Squares for Matrix Parameter Estimation with Application to Adaptive MPC

Abstract
Traditionally, batch least squares (BLS) and recursive least squares (RLS) are used for identification of a vector of parameters that form a linear model. In some situations, however, it is of interest to identify parameters in a matrix structure. In this case, a common approach is to transform the problem into standard vector form using the vectorization (vec) operator and the Kronecker product, known as vec-permutation. However, the use of the Kronecker product introduces extraneous zero terms in the regressor, resulting in unnecessary additional computational and space requirements. This work derives matrix BLS and RLS formulations which, under mild assumptions, minimize the same cost as the vec-permutation approach. This new approach requires less computational complexity and space complexity than vec-permutation in both BLS and RLS identification. It is also shown that persistent excitation guarantees convergence to the true matrix parameters. This method can used to improve computation time in the online identification of multiple-input, multiple-output systems for indirect adaptive model predictive control.
I Introduction
Least squares based identification methods are foundational to systems and control theory, particularly identification, signal processing, and adaptive control [1, 2]. Batch least squares (BLS) and recursive least squares (RLS) are traditionally used to identify a vector of parameters in a linear measurement process [2, 3]. However, it may be of interest to identify parameters in a matrix structure, for example, in adaptive control of multiple-input, multiple-output (MIMO) systems [4, 3]. One approach is to use vec-permutation [5], a method which rewrites the linear measurement process such that the columns of the parameters to be identified are stacked into a vector. This is accomplished using the the vectorization operator and Kronecker product, and is a straightforward solution for various situations [6, 4, 7, 8, 9, 10, 11, 12].
A significant drawback, however, is that the vec-permutation method increases the dimension of the linear measurement process by using the Kronecker product, introducing extraneous zero terms in the regressor (e.g. equation (15) of [4]). This results in increased computational cost and storage requirements. Other methods for identifying parameters in a matrix structure have been proposed including square root filtering [13], multiinnovations [14, 15], and gradient-based methods [16]. However, these methods do not address whether a least squares cost function is globally minimized or the relationship of the method to standard least squares methods.
This work develops a batch and recursive least squares algorithm for identification of matrix parameters which, under mild assumptions, minimizes the same cost function used in the vec-permutation approach. This method provides a $\mathcal{O}(m^{3})$ times improvement in computational cost and a $\mathcal{O}(m^{2})$ times improvement in storage requirements over vec-permutation, where $m\geq 1$ is the number of columns of the identified parameter matrix. We also show how persistent excitation guarantees convergence of the identified matrix parameters to true matrix parameters, which extends established results for identification of vector parameters [17, 18, 19]. Finally, we show how this method can be used to significantly reduce computation time spent on online identification in predictive cost adaptive control (PCAC) [4].
II The Vec-permutation Approach
Consider a measurement process of the form [^1] Note that since the measurement, regressor, and parameters are all matrices, the results of this work can be easily extended to measurements processes of the form $y_{k}=\theta\phi_{k}$ by rewriting as $y_{k}^{\rm T}=\phi_{k}^{\rm T}\theta^{\rm T}$ and identifying parameters $\theta^{\rm T}$. For brevity, we leave the details to the reader.
$y_{k}=\phi_{k}\theta,$ (1)
where $k=0,1,2,\ldots$ is the time step, $y_{k}\in{\mathbb{R}}^{p\times m}$ is the measurement at step $k$, $\phi_{k}\in{\mathbb{R}}^{p\times n}$ is the regressor at step $k$, and $\theta\in{\mathbb{R}}^{n\times m}$ is a matrix of unknown parameters. Parameters $\theta$ can be identified by minimizing the least squares cost function $J_{k}\colon{\mathbb{R}}^{n\times m}\rightarrow{\mathbb{R}}$, defined as
$J_{k}(\hat{\theta})=$ $\sum_{i=0}^{k}\operatorname{vec}(y_{i}-\phi_{i}\hat{\theta})^{\rm T}\bar{\Gamma}_{i}\operatorname{vec}(y_{i}-\phi_{i}\hat{\theta})$
$+\operatorname{vec}(\hat{\theta}-\theta_{0})^{\rm T}\bar{R}\operatorname{vec}(\hat{\theta}-\theta_{0}),$ (2)
where $\operatorname{vec}(\cdot)$ is the column stacking operator, positive-definite $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ is the regularization matrix, $\theta_{0}\in{\mathbb{R}}^{n\times m}$ is an initial estimate of $\theta$, and, for all $k\geq 0$, positive-definite $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ is the measurement weighting matrix.
Using vec-permutation [5], (1) can be rewritten as
$\bar{y}_{k}=\bar{\phi}_{k}\bar{\theta},$ (3)
where $\bar{y}_{k}\in{\mathbb{R}}^{mp}$, $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$, and $\bar{\theta}\in{\mathbb{R}}^{mn}$ are defined
$\bar{y}_{k}$ $\triangleq\operatorname{vec}(y_{k}),$ (4)
$\bar{\phi}_{k}$ $\triangleq(I_{m}\otimes\phi_{k}),$ (5)
$\bar{\theta}$ $\triangleq\operatorname{vec}(\theta),$ (6)
and where $\otimes$ is the Kronecker product. Note, for all $k\geq 0$ and $\hat{\theta}\in{\mathbb{R}}^{n\times m}$ that $\operatorname{vec}(y_{k}-\phi_{k}\hat{\theta})=\bar{y}_{k}-\bar{\phi}_{k}\operatorname{vec}(\hat{\theta})$ and $\operatorname{vec}(\hat{\theta}-\theta_{0})=\operatorname{vec}(\hat{\theta})-\bar{\theta}_{0}$, where $\bar{\theta}_{0}\in{\mathbb{R}}^{mn}$ is defined as
$\bar{\theta}_{0}\triangleq\operatorname{vec}(\theta_{0}).$ (7)
It then follows that, for all $k\geq 0$, the cost function $J_{k}$, given in (II), can be rewritten as
$J_{k}(\hat{\theta})=$ $\sum_{i=0}^{k}(\bar{y}_{i}-\bar{\phi}_{i}\operatorname{vec}(\hat{\theta}))^{\rm T}\bar{\Gamma}_{i}(\bar{y}_{i}-\bar{\phi}_{i}\operatorname{vec}(\hat{\theta}))$
$+(\operatorname{vec}(\hat{\theta})-\bar{\theta}_{0})^{\rm T}\bar{R}(\operatorname{vec}(\hat{\theta})-\bar{\theta}_{0}).$ (8)
Propositions 1 and 2 respectively give the batch and recursive least squares methods using vec-permutation to minimize cost function $J_{k}$.
Proposition 1.
Let $N\geq 0$. For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite. Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (II), has a unique minimizer whose vectorization is given by
$\operatorname{vec}\left(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})\right)=\bar{A}_{N}^{-1}\bar{b}_{N},$ (9)
where
$\bar{A}_{N}$ $\triangleq\bar{R}+\sum_{i=0}^{N}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{\phi}_{i},$ (10)
$\bar{b}_{N}$ $\triangleq\bar{R}\operatorname{vec}(\theta_{0})+\sum_{i=0}^{N}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{y}_{i},$ (11)
and where, for all $0\leq k\leq N$, $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively.
Proof.
This result follows directly from Lemma A.3. ∎
Proposition 2.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{i}\in{\mathbb{R}}^{mp\times mp}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (II), has a unique minimizer, whose vectorization is denoted as
$\bar{\theta}_{k+1}\triangleq\operatorname{vec}\left(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})\right).$ (12)
Moreover, for all $k\geq 0$, $\bar{\theta}_{k+1}$ is given recursively by
$\bar{P}_{k+1}^{-1}$ $=\bar{P}_{k}^{-1}+\bar{\phi}_{k}^{\rm T}\bar{\Gamma}_{k}\bar{\phi}_{k},$ (13)
$\bar{\theta}_{k+1}$ $=\bar{\theta}_{k}+P_{k+1}\bar{\phi}_{k}^{\rm T}\bar{\Gamma}_{k}(\bar{y}_{k}-\bar{\phi}_{k}\bar{\theta}_{k}).$ (14)
where $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively, $\bar{P}_{0}\triangleq\bar{R}^{-1}$, and, for all $k\geq 0$, $\bar{P}_{k}\in{\mathbb{R}}^{mn\times mn}$ is positive definite, and hence nonsingular. Moreover, for all $k\geq 0$, $\bar{P}_{k}$ can be expressed recursively as
$\bar{P}_{k+1}$ $=\bar{P}_{k}-\bar{P}_{k}\bar{\phi}_{k}^{\rm T}(\bar{\Gamma}_{k}^{-1}+\bar{\phi}_{k}\bar{P}_{k}\bar{\phi}_{k}^{\rm T})^{-1}\bar{\phi}_{k}\bar{P}_{k}.$ (15)
Proof.
This result follows directly from Lemma A.3. ∎
An inefficiency with this method is that the Kronecker product in (5) introduces extraneous zero terms in $\bar{\phi}_{k}$ when $m>1$, resulting in a sparse and higher dimension regressor matrix. However, there is no way to simplify the results of Propositions 1 and 2 as the regularization matrix $\bar{R}$ and measurement weighting matrices $\bar{\Gamma}_{k}$ are not necessarily sparse. The computational complexities of BLS and RLS with vec-permutation are shown in Tables I and II respectively.
III Independent Column Weighting
Next, we assume there exist positive-definite $R_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
$\bar{R}=\begin{bmatrix}R_{1}&\cdots&0_{n\times n}\\ \vdots&\ddots&\vdots\\ 0_{n\times n}&\cdots&R_{m}\end{bmatrix}.$ (16)
Furthermore, assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{1,k},\ldots,\Gamma_{m,k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
$\bar{\Gamma}_{k}=\begin{bmatrix}\Gamma_{1,k}&\cdots&0_{p\times p}\\ \vdots&\ddots&\vdots\\ 0_{p\times p}&\cdots&\Gamma_{m,k}\end{bmatrix}.$ (17)
This corresponds to independent weighting of the columns of $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and $\hat{\theta}-\theta_{0}$ in (II). Then, for all $k\geq 0$ and $\hat{\theta}\in{\mathbb{R}}^{n\times m}$, (II) can be rewritten as
$J_{k}(\hat{\theta})=\sum_{j=0}^{m}J_{j,k}(\hat{\theta}_{j}),$ (18)
where, for all $j=1,\ldots,m$, $J_{j,k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is defined as
$J_{j,k}(\hat{\theta}_{j})=$ $\sum_{i=0}^{k}(y_{j,i}-\phi_{i}\hat{\theta}_{j})^{\rm T}\Gamma_{j,i}(y_{j,i}-\phi_{i}\hat{\theta}_{j})$
$+(\hat{\theta}_{j}-\theta_{j,0})^{\rm T}R_{j}(\hat{\theta}_{j}-\theta_{j,0}).$ (19)
where the vectors $y_{1,k},\ldots,y_{m,k}\in{\mathbb{R}}^{p}$, $\theta_{1,0},\ldots,\theta_{m,0}\in{\mathbb{R}}^{n}$, and $\hat{\theta}_{1},\ldots,\hat{\theta}_{m}\in{\mathbb{R}}^{n}$ are the $m$ columns of $y_{k}$, $\theta_{0}$, and $\hat{\theta}$, respectively. In particular,
$y_{k}$ $\triangleq\begin{bmatrix}y_{1,k}&\cdots&y_{m,k}\end{bmatrix},$ (20)
$\theta_{0}$ $\triangleq\begin{bmatrix}\theta_{1,0}&\cdots&\theta_{m,0}\end{bmatrix},$ (21)
$\hat{\theta}$ $\triangleq\begin{bmatrix}\hat{\theta}_{1}&\cdots&\hat{\theta}_{m}\end{bmatrix}.$ (22)
The following Lemma shows that minimizing the cost function $J_{k}$, given by (18), can be done by separately minimizing $J_{j,k}$ for all $j=1,\ldots,m$.
Lemma 1.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer given by
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})$ $=\begin{bmatrix}\operatorname*{arg\,min}\limits_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{1,k}(\hat{\theta}_{1})&\ldots&\operatorname*{arg\,min}\limits_{\hat{\theta}_{m}\in{\mathbb{R}}^{n}}J_{m,k}(\hat{\theta}_{m})\end{bmatrix}$
Proof.
Note that, for all $j=1,\ldots,m$, since $J_{j,k}$ is a function of only $\hat{\theta}_{j}$, it follows from Lemma A.3 that $J_{j,k}$ has a unique minimizer. Then, since $J_{k}(\hat{\theta})=\sum_{j=0}^{m}J_{j,k}(\hat{\theta}_{j})$, it follows from (22) that the Lemma holds. ∎
Propositions 3 and 4 respectively give the independent column weighting batch and recursive least squares methods to minimize cost function $J_{k}$ given by (18).
Proposition 3.
Let $N\geq 0$. For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer given by
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=\begin{bmatrix}A_{1,N}^{-1}b_{1,N}&\cdots&A_{m,N}^{-1}b_{m,N}\end{bmatrix},$ (23)
where, for all $j=1,\ldots,m$,
$A_{j,N}$ $\triangleq R_{j}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{j,i}\phi_{i},$ (24)
$b_{j,N}$ $\triangleq R_{j}\theta_{j,0}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{j,i}y_{j,i},$ (25)
and where, for all $0\leq k\leq N$, $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (20) and $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (21).
Proof.
It follows from Lemma A.3 that, for all $j=1,\ldots,m$, $\operatorname*{arg\,min}_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{j,N}(\hat{\theta}_{j})=A_{j,N}^{-1}b_{j,N}$. Hence, Lemma 1 implies (23). ∎
Proposition 4.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer, denoted as
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix},$ (26)
where, for all $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given by
$P_{j,k+1}^{-1}$ $=P_{j,k}^{-1}+\phi_{k}^{\rm T}\Gamma_{j,k}\phi_{k},$ (27)
$\theta_{j,k+1}$ $=\theta_{j,k}+P_{j,k+1}\phi_{k}^{\rm T}\Gamma_{j,k}(y_{j,k}-\phi_{k}\theta_{j,k}).$ (28)
where $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (20), $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (21), $P_{j,0}\triangleq R_{j}^{-1}$, and, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}\in{\mathbb{R}}^{n\times n}$ is positive definite, and hence nonsingular. Moreover, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}$ can be expressed recursively as
$P_{j,k+1}$ $=P_{j,k}-P_{j,k}\phi_{k}^{\rm T}(\Gamma_{j,k}^{-1}+\phi_{k}P_{j,k}\phi_{k}^{\rm T})^{-1}\phi_{k}P_{j,k}.$ (29)
Proof.
For all $j=1,\ldots,m$, let $\theta_{j,k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{j,k}$. It then follows from Lemma A.3 that, for all $j=1,\ldots,m$, (27) and (28) hold. Finally, Lemma 1 implies (23). ∎
An advantage of independent column weighting versus vec-permutation is that no Kronecker product is used, implying that no sparse matrices are introduced. The computational complexities of BLS and RLS with independent column weighting are shown in Tables I and II respectively. Independent column weighting results in a $\mathcal{O}(m^{2})$ times improvement in computational complexity over vec-permutation for both BLS and RLS as well as a $\mathcal{O}(m)$ times improvement in space complexity.
IV Independent Identical Column Weighting
Finally, we now assume there exist positive-definite $R\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
$\bar{R}=\begin{bmatrix}R&\cdots&0_{n\times n}\\ \vdots&\ddots&\vdots\\ 0_{n\times n}&\cdots&R\end{bmatrix}=I_{m}\otimes R.$ (30)
Furthermore, assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
$\bar{\Gamma}_{k}=\begin{bmatrix}\Gamma_{k}&\cdots&0_{p\times p}\\ \vdots&\ddots&\vdots\\ 0_{p\times p}&\cdots&\Gamma_{k}\end{bmatrix}=I_{m}\otimes\Gamma_{k}.$ (31)
This corresponds to independent and identical weighting of the columns of $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and $\hat{\theta}-\theta_{0}$ in (II). Then, it follows from Lemma A.2 that, for all $k\geq 0$, (II) can be rewritten as
$J_{k}(\hat{\theta})=\operatorname{tr}\bigg{[}$ $\sum_{i=0}^{k}(y_{i}-\phi_{i}\hat{\theta})^{\rm T}\Gamma_{i}(y_{i}-\phi_{i}\hat{\theta})$
$+(\theta-\theta_{0})^{\rm T}R(\theta-\theta_{0})\bigg{]}.$ (32)
Propositions 5 and 6 respectively give the independent identical column weighting batch and recursive least squares methods to minimize cost function $J_{k}$ given by (IV).
Proposition 5.
Let $N\geq 0$. For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (IV), has a unique minimizer given by
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=A_{N}^{-1}b_{N},$ (33)
where
$A_{N}$ $\triangleq R+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i},$ (34)
$b_{N}$ $\triangleq R\theta_{0}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{i}y_{i}.$ (35)
Proof.
Note that (IV) can be rewritten as (18) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$. It then follows from Proposition 3 that $\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=\begin{bmatrix}A_{N}^{-1}b_{1,N}&\cdots&A_{N}^{-1}b_{m,N}\end{bmatrix}=A_{N}^{-1}\begin{bmatrix}b_{1,N}&\cdots&b_{m,N}\end{bmatrix}$, where $A_{N}\in{\mathbb{R}}^{n\times n}$ is defined in (34) and $b_{j,N}$ is defined in (25). Finally, note that $b_{N}=[b_{1,N}\ \cdots\ b_{m,N}]$, thus yielding (33). ∎
Proposition 6.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (IV), has a unique minimizer, denoted as
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\theta_{k+1}.$ (36)
Moreover, for all $k\geq 0$, $\theta_{k+1}\in{\mathbb{R}}^{n\times m}$ is given by
$P_{k+1}^{-1}$ $=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$ (37)
$\theta_{k+1}$ $=\theta_{k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{k}-\phi_{k}\theta_{k}).$ (38)
where $P_{0}\triangleq R^{-1}$ and, for all $k\geq 0$, $P_{k}\in{\mathbb{R}}^{n\times n}$ is positive definite, hence nonsingular. Finally, for all $k\geq 0$, $P_{k+1}$ can be expressed recursively as
$P_{k+1}$ $=P_{k}-P_{k}\phi_{k}^{\rm T}(\Gamma_{k}^{-1}+\phi_{k}P_{k}\phi_{k}^{\rm T})^{-1}\phi_{k}P_{k}.$ (39)
Proof.
Note that (IV) can be rewritten as (18) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$. It then follows from Proposition 4 that, for all $k\geq 0$,
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix},$
where, for all $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given by
$P_{k+1}^{-1}$ $=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$
$\theta_{j,k+1}$ $=\theta_{j,k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{j,k}-\phi_{k}\theta_{j,k}).$
Applying (20) and (22) then implies (37) and (38). ∎
Note that in independent identical column weighting, the same covariance matrix is used for each column of $\theta_{k}$ versus independent column weighting which computes a separate covariance matrix for each column of parameters $\theta_{j,k}$, $j=1,\ldots,m$. The computational complexities of BLS and RLS with independent identical column weighting are shown in Tables I and II respectively. Independent column weighting results in a $\mathcal{O}(m^{3})$ times improvement in computational complexity over vec-permutation as well as a $\mathcal{O}(m^{2})$ times improvement in space complexity over vec-permutation.
Figure 1 shows numerical testing of BLS and RLS with vec-permutation, independent column weighting, and independent identical column weighting. We consider the measurement process (1) with $p=10$, $n=50$, and $1\leq m\leq 20$. For larger values of $m$, we see significantly faster computation time for independent identical column weighting over vec-permutation and independent column weighting.
IV-A Convergence of Matrix RLS
It is well-known that in standard RLS, the parameter estimate vector converges to the vector of true parameters if the sequence of regressors $(\phi_{k})_{k=0}^{\infty}$ is persistently exciting [17, 18, 19]. Theorem 1 extends this result to matrix RLS. The following definition is from page 64 of [1].
Definition 1.
$(\phi_{k})_{k=0}^{\infty}\subset{\mathbb{R}}^{p\times n}$ with weight $(\Gamma_{k})_{k=0}^{\infty}$ is persistently exciting (PE) if
$C\triangleq\lim_{k\rightarrow\infty}\frac{1}{k}\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i}\in{\mathbb{R}}^{n\times n}$ (40)
exists and is positive definite.
Theorem 1.
Let $\theta,{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite. For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$ be given by (1), let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite, and let $P_{k}\in{\mathbb{R}}^{n\times n}$ and $\theta_{k}\in{\mathbb{R}}^{n\times m}$ be given by (37) and (38), respectively. Assume that $(\phi_{k})_{k=0}^{\infty}$ with weight $(\Gamma_{k})_{k=0}^{\infty}$ is PE, and define $C\in{\mathbb{R}}^{n\times n}$ by (40). Then,
$\lim_{k\rightarrow\infty}k(\theta_{k}-\theta)=C^{-1}R(\theta_{0}-\theta).$ (41)
Proof. Note that
$\theta_{k}$ $=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}(R\theta_{0}+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}y_{i})$
$=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}(R\theta_{0}-R\theta+R\theta+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i}\theta)$
$=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}\Big{[}R(\theta_{0}-\theta)+(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})\theta\Big{]}$
$=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}R(\theta_{0}-\theta)+\theta.$
Hence, it follows that
$\lim_{k\rightarrow\infty}k(\theta_{k}-\theta)$ $=\lim_{k\rightarrow\infty}(\frac{1}{k}R+\frac{1}{k}\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}R(\theta_{0}-\theta)$
$=C^{-1}R(\theta_{0}-\theta).$square
TABLE I: Batch identification summary and computational complexities for $N$ measurements
Algorithm Comp. Complexity (No Assumptions)
Comp. Complexity
$(N\gg n\geq p)$
Comp. Complexity
$(N\gg p\geq n)$
Vec-Permutation (9), (10), (11) $\mathcal{O}(\max\{Npnm^{3}\max\{p,n\},n^{3}m^{3}\})$ $\mathcal{O}(Npn^{2}m^{3})$ $\mathcal{O}(Np^{2}nm^{3})$
Indep. Column Weight (23), (24), (25) $\mathcal{O}(\max\{Npnm\max\{p,n\},n^{3}m\})$ $\mathcal{O}(Npn^{2}m)$ $\mathcal{O}(Np^{2}nm)$
Indep. Iden. Column Weight (33), (34), (35) $\mathcal{O}(\max\{Np\max\{n,m\}\max\{p,n\},n^{3}\})$ $\mathcal{O}(Npn\max\{n,m\})$ $\mathcal{O}(Np^{2}\max\{n,m\})$
TABLE II: Recursive identification summary and computational/space complexities per step
Algorithm
$(n\gg p)$
Comp. Complexity
$(n\gg p)$
Algorithm
$(p\geq n)$
Comp. Complexity
$(p\geq n)$
Number of Parameters
in Memory
Vec-Permutation (15), (14) $\mathcal{O}(pn^{2}m^{3})$ (13), (14) $\mathcal{O}(p^{2}nm^{3})$ $n^{2}m^{2}+nm$
Indep. Column Weight (29), (28) $\mathcal{O}(pn^{2}m)$ (27), (28) $\mathcal{O}(p^{2}nm)$ $n^{2}m+nm$
Indep. Iden. Column Weight (39), (38) $\mathcal{O}(pn\max\{n,m\})$ (37), (38) $\mathcal{O}(p^{2}\max\{n,m\})$ $n^{2}+nm$
[…]
V Application to Online Identification for Indirect Adaptive Model Predictive Control
Consider a MIMO input-output system of the form
$y_{k}=-\sum_{i=1}^{\hat{n}}F_{i}y_{k-i}+\sum_{i=0}^{\hat{n}}G_{i}u_{k-i},$ (42)
where $k\geq 0$ is the time step, $\hat{n}$ is the model order, $u_{k}\in{\mathbb{R}}^{m}$ is the control, $y_{k}\in{\mathbb{R}}^{p}$ is the measurement, and $F_{1},\ldots,F_{\hat{n}}\in{\mathbb{R}}^{p\times p}$ and $G_{0},\ldots,G_{\hat{n}}\in{\mathbb{R}}^{p\times m}$ are the system coefficient matrices to be estimated. A model of the form (42) is identified online in the indirect adaptive model predictive control scheme: predictive cost adaptive control (PCAC) [4]. For all $k\geq 0$, the system coefficient matrices are estimated by minimizing the cost function $J_{k}\colon{\mathbb{R}}^{p\times\hat{n}(m+p)+m}\rightarrow{\mathbb{R}}$, defined as
$J_{k}(\hat{\theta})=\sum_{i=0}^{k}z_{i}^{\rm T}(\hat{\theta})z_{i}(\hat{\theta})+\operatorname{vec}(\hat{\theta}-\theta_{0})^{\rm T}\bar{P}_{0}^{-1}\operatorname{vec}(\hat{\theta}-\theta_{0}),$ (43)
where $z_{k}\colon{\mathbb{R}}^{p\times\hat{n}(m+p)+m}\rightarrow{\mathbb{R}}^{p}$ is defined
$z_{k}(\hat{\theta})\triangleq y_{k}+\sum_{i=1}^{\hat{n}}\hat{F}_{i}y_{k-i}-\sum_{i=0}^{\hat{n}}\hat{G}_{i}u_{k-i},$ (44)
$\hat{\theta}\in{\mathbb{R}}^{p\times\hat{n}(m+p)+m}$ are the coefficients to be estimated, defined
$\hat{\theta}\triangleq\begin{bmatrix}\hat{F}_{1}&\cdots&\hat{F}_{\hat{n}}&\hat{G}_{0}&\cdots&\hat{G}_{\hat{n}}\end{bmatrix},$ (45)
and where $\theta_{0}\in{\mathbb{R}}^{p\times\hat{n}(m+p)+m}$ is an initial guess of the coefficients and $\bar{P}_{0}\in{\mathbb{R}}^{[\hat{n}p(m+p)+mp]\times[\hat{n}p(m+p)+mp]}$ is positive definite. Note that, for all $k\geq 0$, $z_{k}(\hat{\theta})$ can be written as
$z_{k}(\hat{\theta})=y_{k}-\hat{\theta}\phi_{k},$ (46)
where $\phi_{k}\in{\mathbb{R}}^{\hat{n}(m+p)+m}$ is defined as
$\phi_{k}\triangleq\begin{bmatrix}-y_{k-1}^{\rm T}&\cdots-y_{k-\hat{n}}^{\rm T}&u_{k}^{\rm T}&\cdots&u_{k-\hat{n}}^{\rm T}\end{bmatrix}^{\rm T}.$ (47)
Further defining $\bar{\phi}_{k}\in{\mathbb{R}}^{p\times\hat{n}p(m+p)+mp}$ as
$\bar{\phi}_{k}\triangleq\phi_{k}^{\rm T}\otimes I_{p},$ (48)
it follows that $z_{k}(\hat{\theta})$ can be written as
$z_{k}(\hat{\theta})=y_{k}-\bar{\phi}_{k}\operatorname{vec}(\hat{\theta})$ (49)
where $\operatorname{vec}(\hat{\theta})\in{\mathbb{R}}^{\hat{n}p(m+p)+mp}$ is the vectorization of $\hat{\theta}$. Using (49), we derive the identification algorithm used in [4].
Proposition 7.
For all $k\geq 0$, let $u_{k}\in{\mathbb{R}}^{m}$, $y_{k}\in{\mathbb{R}}^{p}$. Furthermore, let $\theta_{0}\in{\mathbb{R}}^{p\times\hat{n}(m+p)+m}$ and let $\bar{P}_{0}\in{\mathbb{R}}^{[\hat{n}p(m+p)+mp]\times[\hat{n}p(m+p)+mp]}$ be positive definite. Then, for all $k\geq 0$, $J_{k}$, defined in (43), has a unique global minimizer, denoted
$\theta_{k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{p\times\hat{n}(m+p)+m}}J_{k}(\hat{\theta}),$ (50)
which is given by
$\bar{P}_{k+1}$ $=\bar{P}_{k}-\bar{P}_{k}\bar{\phi}_{k}^{\rm T}(I_{p}+\bar{\phi}_{k}\bar{P}_{k}\bar{\phi}_{k}^{\rm T})^{-1}\bar{\phi}_{k}\bar{P}_{k},$ (51)
$\operatorname{vec}(\theta_{k+1})$ $=\operatorname{vec}(\theta_{k})+\bar{P}_{k+1}\bar{\phi}_{k}^{\rm T}(y_{k}-\bar{\phi}_{k}\operatorname{vec}(\theta_{k})).$ (52)
Proof.
This result follows from Lemma A.3. For further details, see equations (8) through (20) of [4]. ∎
Next, we provide an alternate formulation using matrix RLS.
Proposition 8.
Consider the notation and assumptions of Proposition 7. If there exists $P_{0}\in{\mathbb{R}}^{[\hat{n}(m+p)+m]\times[\hat{n}(m+p)+m]}$ such that $\bar{P_{0}}=P_{0}\otimes I_{p}$, then, for all $k\geq 0$, $\theta_{k+1}\in{\mathbb{R}}^{p\times\hat{n}(m+p)+m}$ is given by
$P_{k+1}$ $=P_{k}-\frac{P_{k}\phi_{k}\phi_{k}^{\rm T}P_{k}}{1+\phi_{k}^{\rm T}P_{k}\phi_{k}},$ (53)
$\theta_{k+1}$ $=\theta_{k}+(y_{k}-\theta_{k}\phi_{k})\phi_{k}^{\rm T}P_{k+1}.$ (54)
Proof.
Note that, for all $k\geq 0$, $z_{k}(\hat{\theta})^{\rm T}z_{k}(\hat{\theta})=\operatorname{tr}(z_{k}(\hat{\theta})^{\rm T}z_{k}(\hat{\theta}))=\operatorname{tr}(z_{k}(\hat{\theta})z_{k}(\hat{\theta})^{\rm T})$. It then follows from (57) of Lemma A.2 that, for all $k\geq 0$, (43) can be written as
$J_{k}(\hat{\theta})=\operatorname{tr}\Big{[}\sum_{i=0}^{k}z_{i}(\hat{\theta})z_{i}(\hat{\theta})^{\rm T}+(\hat{\theta}-\theta_{0}){P}_{0}^{-1}(\hat{\theta}-\theta_{0})^{\rm T}\Big{]}.$
Proposition 5 then implies that, for all $k\geq 0$, (53) and $\theta_{k+1}^{\rm T}=\theta_{k}^{\rm T}+P_{k+1}\phi_{k}(y_{k}^{\rm T}-\phi_{k}^{\rm T}\theta_{k}^{\rm T})$ hold. Taking the transpose then yields (54). ∎
Example 1. This example is from [10] and uses PCAC for the control of a flexible structure under harmonic and broadband disturbances. Consider the 4-bay truss show in Figure 2 made of flexible truss elements with unknown mass and stiffness. Two actuators are placed at nodes 3 and 4 with control authority in the $x$-direction and $x$-direction displacement sensors are placed at nodes 5, 6, 7, and 8. The objective is to use PCAC to suppress the effects of harmonic and broadband disturbances without prior knowledge of the truss dynamics. See [10] for further details.
[…]
Figure 2: Flexible truss structure from [10] with nodes labeled.
TABLE III: Truss ex. computation time per step: mean $\pm$ std. deviation
Vec-Permutation Matrix RLS Change
[…]
VI Conclusions
This work derives batch and recursive least squares algorithms for the identification of matrix parameters. Under the assumption of independent, identical column weighting, these methods minimize the same cost function as the vec-permutation approach. It is also shown how, under persistent excitation, convergence guarantees can be extended from the vector case to the matrix case. This approach can be used fast online identification of MIMO systems which is critical in indirect adaptive model predictive control.

### v2

@@
-Efficient Batch and Recursive Least Squares for Matrix Parameter Estimation with Application to Adaptive MPC
+Efficient Batch and Recursive Least Squares for Matrix Parameter Estimation
 Abstract
-Traditionally, batch least squares (BLS) and recursive least squares (RLS) are used for identification of a vector of parameters that form a linear model.
-In some situations, however, it is of interest to identify parameters in a matrix structure.
+Traditionally, batch least squares (BLS) and recursive least squares (RLS) are used for identification of a vector of parameters which form a linear model.
+In some situations however, it is of interest to identify parameters in a matrix structure.
 In this case, a common approach is to transform the problem into standard vector form using the vectorization (vec) operator and the Kronecker product, known as vec-permutation.
@@
 This method can used to improve computation time in the online identification of multiple-input, multiple-output systems for indirect adaptive model predictive control.
-I Introduction
+{IEEEkeywords}
+Identification, Modeling, Adaptive Systems, MIMO Systems
+1 Introduction
 Least squares based identification methods are foundational to systems and control theory, particularly identification, signal processing, and adaptive control [1, 2].
@@
 This results in increased computational cost and storage requirements.
-Other methods for identifying parameters in a matrix structure have been proposed including square root filtering [13], multiinnovations [14, 15], and gradient-based methods [16].
-However, these methods do not address whether a least squares cost function is globally minimized or the relationship of the method to standard least squares methods.
-This work develops a batch and recursive least squares algorithm for identification of matrix parameters which, under mild assumptions, minimizes the same cost function used in the vec-permutation approach.
-This method provides a $\mathcal{O}(m^{3})$ times improvement in computational cost and a $\mathcal{O}(m^{2})$ times improvement in storage requirements over vec-permutation, where $m\geq 1$ is the number of columns of the identified parameter matrix.
-We also show how persistent excitation guarantees convergence of the identified matrix parameters to true matrix parameters, which extends established results for identification of vector parameters [17, 18, 19].
+Another approach is to apply standard least squares methods to separately identify the columns of the matrix of parameters [2, p.
+102] but this method does not address what cost function is being minimized or the relationship to the vec-permutation approach.
+Other related methods including square root filtering [13], multiinnovations [14], and gradient-based methods [15] also do not address whether a least squares cost function is globally minimized or the relationship to standard least squares methods.
+This work derives a batch and recursive least squares algorithm for identification of matrix parameters which, under the assumption of independent residual error and parameter column weighting, minimizes the same cost function used in the vec-permutation approach.
+This method provides an $\mathcal{O}(m^{3})$ times improvement in computational complexity and an $\mathcal{O}(m^{2})$ times improvement in storage requirements over vec-permutation, where $m\geq 1$ is the number of columns of the identified parameter matrix.
+We also show how persistent excitation guarantees convergence of the identified matrix parameters to true matrix parameters, which extends established results for identification of vector parameters [16, 17].
+However, we show this improvement in computational complexity may come at the cost of performance if the columns of measurement noise are highly correlated.
 Finally, we show how this method can be used to significantly reduce computation time spent on online identification in predictive cost adaptive control (PCAC) [4].
-II The Vec-permutation Approach
+2 Vec-permutation Least Squares
 Consider a measurement process of the form [^1] Note that since the measurement, regressor, and parameters are all matrices, the results of this work can be easily extended to measurements processes of the form $y_{k}=\theta\phi_{k}$ by rewriting as $y_{k}^{\rm T}=\phi_{k}^{\rm T}\theta^{\rm T}$ and identifying parameters $\theta^{\rm T}$.
@@
 $+\operatorname{vec}(\hat{\theta}-\theta_{0})^{\rm T}\bar{R}\operatorname{vec}(\hat{\theta}-\theta_{0}),$ (2)
-where $\operatorname{vec}(\cdot)$ is the column stacking operator, positive-definite $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ is the regularization matrix, $\theta_{0}\in{\mathbb{R}}^{n\times m}$ is an initial estimate of $\theta$, and, for all $k\geq 0$, positive-definite $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ is the measurement weighting matrix.
+where $\operatorname{vec}(\cdot)$ is the column stacking operator, positive-definite (and thus, by definition, symmetric) $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ is the regularization matrix, $\theta_{0}\in{\mathbb{R}}^{n\times m}$ is an initial estimate of $\theta$, and, for all $k\geq 0$, positive-definite $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ is the weighting matrix.
+If no estimate of the parameters $\theta$ is known, common choices in practice for $\theta_{0}$ are $\theta_{0}=0$ or randomly sampled values.
+Note that while the regularization term $\operatorname{vec}(\hat{\theta}-\theta_{0})^{\rm T}\bar{R}\operatorname{vec}(\hat{\theta}-\theta_{0})$ results in a biased estimate, this term reduces variance and guarantees that the least squares cost (2) has a unique global minimizer which is useful when few data have been collected [17][18, sec.
+7.3].
 Using vec-permutation [5], (1) can be rewritten as
@@
 $\bar{y}_{k}$ $\triangleq\operatorname{vec}(y_{k}),$ (4)
-$\bar{\phi}_{k}$ $\triangleq(I_{m}\otimes\phi_{k}),$ (5)
+$\bar{\phi}_{k}$ $\triangleq I_{m}\otimes\phi_{k},$ (5)
 $\bar{\theta}$ $\triangleq\operatorname{vec}(\theta),$ (6)
@@
 $\bar{\theta}_{0}\triangleq\operatorname{vec}(\theta_{0}).$ (7)
-It then follows that, for all $k\geq 0$, the cost function $J_{k}$, given in (II), can be rewritten as
+It then follows that, for all $k\geq 0$, the cost function $J_{k}$, given in (2), can be rewritten as
 $J_{k}(\hat{\theta})=$ $\sum_{i=0}^{k}(\bar{y}_{i}-\bar{\phi}_{i}\operatorname{vec}(\hat{\theta}))^{\rm T}\bar{\Gamma}_{i}(\bar{y}_{i}-\bar{\phi}_{i}\operatorname{vec}(\hat{\theta}))$
 $+(\operatorname{vec}(\hat{\theta})-\bar{\theta}_{0})^{\rm T}\bar{R}(\operatorname{vec}(\hat{\theta})-\bar{\theta}_{0}).$ (8)
-Propositions 1 and 2 respectively give the batch and recursive least squares methods using vec-permutation to minimize cost function $J_{k}$.
+Propositions 1 gives the vec-permutation approach to minimize cost function $J_{k}$.
 Proposition 1.
-Let $N\geq 0$.
-For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ be positive definite.
+For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ be positive definite.
 Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite.
-Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (II), has a unique minimizer whose vectorization is given by
-$\operatorname{vec}\left(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})\right)=\bar{A}_{N}^{-1}\bar{b}_{N},$ (9)
+Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (2), has a unique minimizer, whose vectorization is denoted as $\bar{\theta}_{k+1}\triangleq\operatorname{vec}(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta}))$, which is given by
+$\bar{\theta}_{k+1}=\bar{A}_{k}^{-1}\bar{b}_{k},$ (9)
 where
-$\bar{A}_{N}$ $\triangleq\bar{R}+\sum_{i=0}^{N}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{\phi}_{i},$ (10)
-$\bar{b}_{N}$ $\triangleq\bar{R}\operatorname{vec}(\theta_{0})+\sum_{i=0}^{N}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{y}_{i},$ (11)
-and where, for all $0\leq k\leq N$, $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively.
-Proof.
-This result follows directly from Lemma A.3. ∎
-Proposition 2.
-For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{i}\in{\mathbb{R}}^{mp\times mp}$ be positive definite.
-Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite.
-Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (II), has a unique minimizer, whose vectorization is denoted as
-$\bar{\theta}_{k+1}\triangleq\operatorname{vec}\left(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})\right).$ (12)
-Moreover, for all $k\geq 0$, $\bar{\theta}_{k+1}$ is given recursively by
+$\bar{A}_{k}$ $\triangleq\bar{R}+\sum_{i=0}^{k}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{\phi}_{i},$ (10)
+$\bar{b}_{k}$ $\triangleq\bar{R}\operatorname{vec}(\theta_{0})+\sum_{i=0}^{k}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{y}_{i},$ (11)
+and where, for all $k\geq 0$, $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively.
+Furthermore, for all $k\geq 0$, $\bar{\theta}_{k+1}$ is given recursively by
 $\bar{P}_{k+1}^{-1}$ $=\bar{P}_{k}^{-1}+\bar{\phi}_{k}^{\rm T}\bar{\Gamma}_{k}\bar{\phi}_{k},$ (12)
 $\bar{\theta}_{k+1}$ $=\bar{\theta}_{k}+P_{k+1}\bar{\phi}_{k}^{\rm T}\bar{\Gamma}_{k}(\bar{y}_{k}-\bar{\phi}_{k}\bar{\theta}_{k}).$ (13)
-where $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively, $\bar{P}_{0}\triangleq\bar{R}^{-1}$, and, for all $k\geq 0$, $\bar{P}_{k}\in{\mathbb{R}}^{mn\times mn}$ is positive definite, and hence nonsingular.
+where $\bar{P}_{0}\triangleq\bar{R}^{-1}$, and, for all $k\geq 0$, $\bar{P}_{k}\in{\mathbb{R}}^{mn\times mn}$ is positive definite, and hence nonsingular.
 Moreover, for all $k\geq 0$, $\bar{P}_{k}$ can be expressed recursively as
@@
 Proof.
-This result follows directly from Lemma A.3. ∎
-An inefficiency with this method is that the Kronecker product in (5) introduces extraneous zero terms in $\bar{\phi}_{k}$ when $m>1$, resulting in a sparse and higher dimension regressor matrix.
-However, there is no way to simplify the results of Propositions 1 and 2 as the regularization matrix $\bar{R}$ and measurement weighting matrices $\bar{\Gamma}_{k}$ are not necessarily sparse.
-The computational complexities of BLS and RLS with vec-permutation are shown in Tables I and II respectively.
-III Independent Column Weighting
-Next, we assume there exist positive-definite $R_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
-$\bar{R}=\begin{bmatrix}R_{1}&\cdots&0_{n\times n}\\ \vdots&\ddots&\vdots\\ 0_{n\times n}&\cdots&R_{m}\end{bmatrix}.$ (16)
-Furthermore, assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{1,k},\ldots,\Gamma_{m,k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
-$\bar{\Gamma}_{k}=\begin{bmatrix}\Gamma_{1,k}&\cdots&0_{p\times p}\\ \vdots&\ddots&\vdots\\ 0_{p\times p}&\cdots&\Gamma_{m,k}\end{bmatrix}.$ (17)
-This corresponds to independent weighting of the columns of $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and $\hat{\theta}-\theta_{0}$ in (II).
-Then, for all $k\geq 0$ and $\hat{\theta}\in{\mathbb{R}}^{n\times m}$, (II) can be rewritten as
+Since $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (2), is a standard least squares cost with vector parameters, this result follows directly from [3]. ∎
+Equations (9) through (13) give the batch least squares solution using vec-permutation while (12) through (14) give the recursive least squares solution using vec-permutation.
+An inefficiency with this method is that the Kronecker product in (5) introduces extraneous zero terms in $\bar{\phi}_{k}$ when $m>1$, resulting in a sparse and higher dimensional regressor matrix.
+However, the results of Proposition 1 cannot be simplified since the regularization matrix $\bar{R}$ and weighting matrices $\bar{\Gamma}_{k}$ are not necessarily sparse.
+3 Column-by-Column Least Squares
+To simplify the vec-permutation approach, we make the assumption that there exist positive-definite $R_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
+$\bar{R}=\operatorname{diag}(R_{1},\ldots,R_{m}).$ (15)
+Furthermore, we assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{1,k},\ldots,\Gamma_{m,k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
+$\bar{\Gamma}_{k}=\operatorname{diag}(\Gamma_{1,k},\ldots,\Gamma_{m,k}).$ (16)
+This corresponds to independent weighting of the columns of the residual error, $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and of the regularization term, $\hat{\theta}-\theta_{0}$, in (2).
+Then, for all $k\geq 0$ and $\hat{\theta}\in{\mathbb{R}}^{n\times m}$, (2) can be rewritten as
 $J_{k}(\hat{\theta})=\sum_{j=0}^{m}J_{j,k}(\hat{\theta}_{j}),$ (17)
@@
 $\hat{\theta}$ $\triangleq\begin{bmatrix}\hat{\theta}_{1}&\cdots&\hat{\theta}_{m}\end{bmatrix}.$ (21)
-The following Lemma shows that minimizing the cost function $J_{k}$, given by (18), can be done by separately minimizing $J_{j,k}$ for all $j=1,\ldots,m$.
-Lemma 1.
-For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite.
-Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite.
-Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer given by
-$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})$ $=\begin{bmatrix}\operatorname*{arg\,min}\limits_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{1,k}(\hat{\theta}_{1})&\ldots&\operatorname*{arg\,min}\limits_{\hat{\theta}_{m}\in{\mathbb{R}}^{n}}J_{m,k}(\hat{\theta}_{m})\end{bmatrix}$
-Proof.
-Note that, for all $j=1,\ldots,m$, since $J_{j,k}$ is a function of only $\hat{\theta}_{j}$, it follows from Lemma A.3 that $J_{j,k}$ has a unique minimizer.
-Then, since $J_{k}(\hat{\theta})=\sum_{j=0}^{m}J_{j,k}(\hat{\theta}_{j})$, it follows from (22) that the Lemma holds.
-∎
-Propositions 3 and 4 respectively give the independent column weighting batch and recursive least squares methods to minimize cost function $J_{k}$ given by (18).
-Proposition 3.
-Let $N\geq 0$.
-For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite.
-Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite.
-Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer given by
-$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=\begin{bmatrix}A_{1,N}^{-1}b_{1,N}&\cdots&A_{m,N}^{-1}b_{m,N}\end{bmatrix},$ (23)
-where, for all $j=1,\ldots,m$,
-$A_{j,N}$ $\triangleq R_{j}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{j,i}\phi_{i},$ (24)
-$b_{j,N}$ $\triangleq R_{j}\theta_{j,0}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{j,i}y_{j,i},$ (25)
-and where, for all $0\leq k\leq N$, $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (20) and $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (21).
-Proof.
-It follows from Lemma A.3 that, for all $j=1,\ldots,m$, $\operatorname*{arg\,min}_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{j,N}(\hat{\theta}_{j})=A_{j,N}^{-1}b_{j,N}$. Hence, Lemma 1 implies (23). ∎
+Propositions 2 show that, under assumptions (15) and (16), the cost function $J_{k}$, given by (17), can be minimized by separately updating the columns of the parameter estimate.
+We call this the column-by-column approach.
 Proposition 2.
@@
 Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite.
-Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer, denoted as
-$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix},$ (26)
-where, for all $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given by
+Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (17), has a unique minimizer, whose columns are denoted as
+$\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix}\triangleq\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta}),$ (22)
+which, for all $j=1,\ldots,m$, are given by
+$\theta_{j,k+1}=A_{j,k}^{-1}b_{j,k},$ (23)
+where
+$A_{j,k}$ $\triangleq R_{j}+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{j,i}\phi_{i},$ (24)
+$b_{j,k}$ $\triangleq R_{j}\theta_{j,0}+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{j,i}y_{j,i},$ (25)
+and where, for all $k\geq 0$, $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (19) and $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (20).
+Furthermore, for all $k\geq 0$ and $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given recursively by
 $P_{j,k+1}^{-1}$ $=P_{j,k}^{-1}+\phi_{k}^{\rm T}\Gamma_{j,k}\phi_{k},$ (26)
 $\theta_{j,k+1}$ $=\theta_{j,k}+P_{j,k+1}\phi_{k}^{\rm T}\Gamma_{j,k}(y_{j,k}-\phi_{k}\theta_{j,k}).$ (27)
-where $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (20), $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (21), $P_{j,0}\triangleq R_{j}^{-1}$, and, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}\in{\mathbb{R}}^{n\times n}$ is positive definite, and hence nonsingular.
+where $P_{j,0}\triangleq R_{j}^{-1}$, and, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}\in{\mathbb{R}}^{n\times n}$ is positive definite, and hence nonsingular.
 Moreover, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}$ can be expressed recursively as
@@
 Proof.
-For all $j=1,\ldots,m$, let $\theta_{j,k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{j,k}$.
-It then follows from Lemma A.3 that, for all $j=1,\ldots,m$, (27) and (28) hold.
-Finally, Lemma 1 implies (23).
-∎
-An advantage of independent column weighting versus vec-permutation is that no Kronecker product is used, implying that no sparse matrices are introduced.
-The computational complexities of BLS and RLS with independent column weighting are shown in Tables I and II respectively.
-Independent column weighting results in a $\mathcal{O}(m^{2})$ times improvement in computational complexity over vec-permutation for both BLS and RLS as well as a $\mathcal{O}(m)$ times improvement in space complexity.
-IV Independent Identical Column Weighting
-Finally, we now assume there exist positive-definite $R\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
-$\bar{R}=\begin{bmatrix}R&\cdots&0_{n\times n}\\ \vdots&\ddots&\vdots\\ 0_{n\times n}&\cdots&R\end{bmatrix}=I_{m}\otimes R.$ (30)
-Furthermore, assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
-$\bar{\Gamma}_{k}=\begin{bmatrix}\Gamma_{k}&\cdots&0_{p\times p}\\ \vdots&\ddots&\vdots\\ 0_{p\times p}&\cdots&\Gamma_{k}\end{bmatrix}=I_{m}\otimes\Gamma_{k}.$ (31)
-This corresponds to independent and identical weighting of the columns of $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and $\hat{\theta}-\theta_{0}$ in (II).
-Then, it follows from Lemma A.2 that, for all $k\geq 0$, (II) can be rewritten as
+Note that, for all $j=1,\ldots,m$, $J_{j,k}$ is a function of only $\hat{\theta}_{j}$ and is a standard least squares cost with vector parameters.
+It then follows from [3] that, for all $j=1,\ldots,m$, $J_{j,k}$ has a unique minimizer, denoted as $\theta_{j,k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}_{j}\in{\mathbb{R}}^{n}}J_{j,N}(\hat{\theta}_{j})$, which is given by (23).
+It also follows from [3] that (24) through (28) hold.
+Finally, it follows from (17) and (21) that
+$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})$ $=\begin{bmatrix}\operatorname*{arg\,min}\limits_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{1,k}(\hat{\theta}_{1})&\ldots&\operatorname*{arg\,min}\limits_{\hat{\theta}_{m}\in{\mathbb{R}}^{n}}J_{m,k}(\hat{\theta}_{m})\end{bmatrix}$
+and (22) follows. ∎
+An advantage of the column-by-column approach versus vec-permutation is that no Kronecker product is used, implying that no sparse matrices are introduced.
+Note that the column-by-column approach can also be derived by applying standard least squares methods to the $m$ columns of (1) [2, p.
+102].
+However, our derivation further shows the connection to vec-permutation and how the column-by-column approach implicitly implies independent weighting to the columns of the residual error and parameter regularization.
+4 Matrix Update Least Squares
+Finally, we make the stronger assumption that there exists positive-definite $R\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
+$\bar{R}=\operatorname{diag}(R,\ldots,R)=I_{m}\otimes R.$ (29)
+Furthermore, we assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
+$\bar{\Gamma}_{k}=\operatorname{diag}(\Gamma_{k},\ldots,\Gamma_{k})=I_{m}\otimes\Gamma_{k}.$ (30)
+This corresponds to independent and identical weighting of the columns of the residual error, $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and of the regularization term, $\hat{\theta}-\theta_{0}$, in (2).
+Then, it follows from Lemma A.1 that, for all $k\geq 0$, (2) can be rewritten as
 $J_{k}(\hat{\theta})=\operatorname{tr}\bigg{[}$ $\sum_{i=0}^{k}(y_{i}-\phi_{i}\hat{\theta})^{\rm T}\Gamma_{i}(y_{i}-\phi_{i}\hat{\theta})$
 $+(\theta-\theta_{0})^{\rm T}R(\theta-\theta_{0})\bigg{]}.$ (31)
-Propositions 5 and 6 respectively give the independent identical column weighting batch and recursive least squares methods to minimize cost function $J_{k}$ given by (IV).
-Proposition 5.
-Let $N\geq 0$.
-For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite.
-Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite.
-Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (IV), has a unique minimizer given by
-$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=A_{N}^{-1}b_{N},$ (33)
-where
-$A_{N}$ $\triangleq R+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i},$ (34)
-$b_{N}$ $\triangleq R\theta_{0}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{i}y_{i}.$ (35)
-Proof.
-Note that (IV) can be rewritten as (18) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$.
-It then follows from Proposition 3 that $\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=\begin{bmatrix}A_{N}^{-1}b_{1,N}&\cdots&A_{N}^{-1}b_{m,N}\end{bmatrix}=A_{N}^{-1}\begin{bmatrix}b_{1,N}&\cdots&b_{m,N}\end{bmatrix}$, where $A_{N}\in{\mathbb{R}}^{n\times n}$ is defined in (34) and $b_{j,N}$ is defined in (25).
-Finally, note that $b_{N}=[b_{1,N}\ \cdots\ b_{m,N}]$, thus yielding (33).
-∎
+Propositions 3 show that, under assumptions (29) and (30), the cost function $J_{k}$, given by (4), can be minimized as a single matrix equation. We call this the matrix update approach.
 Proposition 3.
@@
 Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite.
-Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (IV), has a unique minimizer, denoted as
-$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\theta_{k+1}.$ (36)
-Moreover, for all $k\geq 0$, $\theta_{k+1}\in{\mathbb{R}}^{n\times m}$ is given by
+Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (4), has a unique minimizer, denoted as $\theta_{k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})$, which is given by
+$\theta_{k+1}=A_{k}^{-1}b_{k},$ (32)
+where
+$A_{k}$ $\triangleq R+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i},$ (33)
+$b_{k}$ $\triangleq R\theta_{0}+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{i}y_{i}.$ (34)
+Furthermore, for all $k\geq 0$, $\theta_{k+1}\in{\mathbb{R}}^{n\times m}$ is given recursively by
 $P_{k+1}^{-1}$ $=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$ (35)
 $\theta_{k+1}$ $=\theta_{k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{k}-\phi_{k}\theta_{k}).$ (36)
-where $P_{0}\triangleq R^{-1}$ and, for all $k\geq 0$, $P_{k}\in{\mathbb{R}}^{n\times n}$ is positive definite, hence nonsingular. Finally, for all $k\geq 0$, $P_{k+1}$ can be expressed recursively as
+where $P_{0}\triangleq R^{-1}$ and, for all $k\geq 0$, $P_{k}\in{\mathbb{R}}^{n\times n}$ is positive definite, hence nonsingular.
+Moreover, for all $k\geq 0$, $P_{k+1}$ can be expressed recursively as
 $P_{k+1}$ $=P_{k}-P_{k}\phi_{k}^{\rm T}(\Gamma_{k}^{-1}+\phi_{k}P_{k}\phi_{k}^{\rm T})^{-1}\phi_{k}P_{k}.$ (37)
 Proof.
-Note that (IV) can be rewritten as (18) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$. It then follows from Proposition 4 that, for all $k\geq 0$,
-$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix},$
-where, for all $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given by
-$P_{k+1}^{-1}$ $=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$
-$\theta_{j,k+1}$ $=\theta_{j,k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{j,k}-\phi_{k}\theta_{j,k}).$
-Applying (20) and (22) then implies (37) and (38). ∎
-Note that in independent identical column weighting, the same covariance matrix is used for each column of $\theta_{k}$ versus independent column weighting which computes a separate covariance matrix for each column of parameters $\theta_{j,k}$, $j=1,\ldots,m$.
-The computational complexities of BLS and RLS with independent identical column weighting are shown in Tables I and II respectively.
-Independent column weighting results in a $\mathcal{O}(m^{3})$ times improvement in computational complexity over vec-permutation as well as a $\mathcal{O}(m^{2})$ times improvement in space complexity over vec-permutation.
-Figure 1 shows numerical testing of BLS and RLS with vec-permutation, independent column weighting, and independent identical column weighting.
-We consider the measurement process (1) with $p=10$, $n=50$, and $1\leq m\leq 20$.
-For larger values of $m$, we see significantly faster computation time for independent identical column weighting over vec-permutation and independent column weighting.
-IV-A Convergence of Matrix RLS
-It is well-known that in standard RLS, the parameter estimate vector converges to the vector of true parameters if the sequence of regressors $(\phi_{k})_{k=0}^{\infty}$ is persistently exciting [17, 18, 19].
+Note that (4) can be rewritten as (17) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$.
+It then follows from Proposition 2 that $J_{k}$ has a unique minimizer given by $\theta_{k+1}=\begin{bmatrix}A_{k}^{-1}b_{1,k}&\cdots&A_{k}^{-1}b_{m,k}\end{bmatrix}=A_{k}^{-1}\begin{bmatrix}b_{1,k}&\cdots&b_{m,k}\end{bmatrix}$, where $A_{k}\in{\mathbb{R}}^{n\times n}$ is defined in (33) and $b_{1,k},\ldots,b_{m,k}$ are defined in (25).
+Finally, note that $b_{k}=[b_{1,k}\ \cdots\ b_{m,k}]$, yielding (32).
+Moreover, it also follows from Proposition 2 that, for all $k\geq 0$, $P_{k+1}^{-1}=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$ and $\theta_{j,k+1}=\theta_{j,k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{j,k}-\phi_{k}\theta_{j,k})$, where, for all $j=1,\ldots,m$, $\theta_{j,k}\in{\mathbb{R}}^{n}$ and $y_{j,k}\in{\mathbb{R}}^{p}$ are the $j^{\rm th}$ columns of $\theta_{k}$ and $y_{k}$, respectively.
+Combining columns yields (36) and applying matrix inversion lemma to (35) yields (37).
+∎
+Similarly, while the matrix update approach can be derived by applying standard least squares methods to the $m$ columns of (1) and combining columns [2, p.
+103], our derivation shows how the matrix update approach implicitly implies independent and identical weighting to the columns of the residual error and parameter regularization.
+4.1 Convergence of Matrix Update RLS
+It is well-known that in standard RLS, the parameter estimate vector converges to the vector of true parameters if the sequence of regressors $(\phi_{k})_{k=0}^{\infty}$ is persistently exciting [16, 17].
 Theorem 1 extends this result to matrix RLS.
-The following definition is from page 64 of [1].
+To begin, we extend the definition of persistent excitation (PE) from page 64 of [1] to the case of matrix regressors and nonuniform weight.222 Note that [1] considers a measurement process $y_{k}=\phi_{k}^{\rm T}\theta$ while we consider $y_{k}=\phi_{k}\theta$.
+As such, PE is defined in [1] using the limit $\lim_{k\rightarrow\infty}\frac{1}{k}\sum_{i=0}^{k-1}\phi_{i}\phi_{i}^{\rm T}$ whereas we consider $\lim_{k\rightarrow\infty}\frac{1}{k}\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\phi_{i}$.
 Definition 1.
@@
 $\theta_{k}$ $=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}(R\theta_{0}+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}y_{i})$
-$=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}(R\theta_{0}-R\theta+R\theta+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i}\theta)$
 $=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}\Big{[}R(\theta_{0}-\theta)+(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})\theta\Big{]}$
@@
 $=C^{-1}R(\theta_{0}-\theta).$square
-TABLE I: Batch identification summary and computational complexities for $N$ measurements
+Table 1: Batch least squares summary and computational complexities for $N$ measurements
 Algorithm Comp. Complexity (No Assumptions)
@@
 Vec-Permutation (9), (10), (11) $\mathcal{O}(\max\{Npnm^{3}\max\{p,n\},n^{3}m^{3}\})$ $\mathcal{O}(Npn^{2}m^{3})$ $\mathcal{O}(Np^{2}nm^{3})$
-Indep. Column Weight (23), (24), (25) $\mathcal{O}(\max\{Npnm\max\{p,n\},n^{3}m\})$ $\mathcal{O}(Npn^{2}m)$ $\mathcal{O}(Np^{2}nm)$
-Indep. Iden. Column Weight (33), (34), (35) $\mathcal{O}(\max\{Np\max\{n,m\}\max\{p,n\},n^{3}\})$ $\mathcal{O}(Npn\max\{n,m\})$ $\mathcal{O}(Np^{2}\max\{n,m\})$
-TABLE II: Recursive identification summary and computational/space complexities per step
+Column-by-Column (23), (24), (25) $\mathcal{O}(\max\{Npnm\max\{p,n\},n^{3}m\})$ $\mathcal{O}(Npn^{2}m)$ $\mathcal{O}(Np^{2}nm)$
+Matrix Update (32), (33), (34) $\mathcal{O}(\max\{Np\max\{n,m\}\max\{p,n\},n^{3}\})$ $\mathcal{O}(Npn\max\{n,m\})$ $\mathcal{O}(Np^{2}\max\{n,m\})$
+Table 2: Recursive least squares summary and computational/space complexities per step
 Algorithm
@@
 Vec-Permutation (14), (13) $\mathcal{O}(pn^{2}m^{3})$ (12), (13) $\mathcal{O}(p^{2}nm^{3})$ $n^{2}m^{2}+nm$
-Indep. Column Weight (29), (28) $\mathcal{O}(pn^{2}m)$ (27), (28) $\mathcal{O}(p^{2}nm)$ $n^{2}m+nm$
-Indep. Iden. Column Weight (39), (38) $\mathcal{O}(pn\max\{n,m\})$ (37), (38) $\mathcal{O}(p^{2}\max\{n,m\})$ $n^{2}+nm$
+Column-by-Column (28), (27) $\mathcal{O}(pn^{2}m)$ (26), (27) $\mathcal{O}(p^{2}nm)$ $n^{2}m+nm$
+Matrix Update (37), (36) $\mathcal{O}(pn\max\{n,m\})$ (35), (36) $\mathcal{O}(p^{2}\max\{n,m\})$ $n^{2}+nm$
 […]
-V Application to Online Identification for Indirect Adaptive Model Predictive Control
-Consider a MIMO input-output system of the form
+5 Computational Complexity and Performance Tradeoff
+Next, we study the computational complexities of vec-permutation, column-by-column, and matrix update least squares.
+For simplicity, we assume it takes ${\mathcal{O}}(nmp)$ arithmetic operations to multiply an $(n\times m)$ matrix with a $(m\times p)$ matrix, ${\mathcal{O}}(n^{3})$ arithmetic operations to invert an $(n\times n)$ matrix, and ${\mathcal{O}}(nm)$ arithmetic operations to add two $(n\times m)$ matrices.
+The computational complexities of the batch and recursive least squares methods are shown in Tables 1 and 2 respectively.
+Note that for the RLS methods, columns 2 and 4 of Table 2 show the most efficient implementation depending on the dimensions $n$ and $p$.
+Note that for both BLS and RLS, column-by-column and matrix update result in an $\mathcal{O}(m^{2})$ and $\mathcal{O}(m^{3})$ times improvement in computational complexity over vec-permutation, respectively.
+Moreover, for RLS, column-by-column and matrix update result in an $\mathcal{O}(m)$ and $\mathcal{O}(m^{2})$ times improvement in space complexity over vec-permutation, respectively.
+Next, Figure 1 shows numerical testing of BLS and RLS with vec-permutation, column-by-column, and matrix update.
+We consider the measurement process (1) with $p=10$, $n=50$, and $1\leq m\leq 20$.
+For larger values of $m$, we see significantly faster computation time for matrix update over vec-permutation and column-by-column.
+While matrix update least squares offers a significant improvement in computational cost for large values of $m$, the following example shows that there may be a sacrifice in performance if there is prior knowledge that measurement noise has highly correlated columns.
+While not shown in this example, there may also be performance sacrifice if there is prior knowledge that columns of the parameters are highly correlated.
+Example 5.1.
+Consider the measurement process (1) with $p=2$, $m=2$, and $n=100$.
+We consider $10$ independent trials where, for each trial, the two columns of the parameters $\theta$ are i.i.d.
+sampled from the Gaussian distribution $\mathcal{N}(0_{n\times 1},I_{n})$.
+For each trial, for all $k\geq 0$, the two rows of the regressor, ${\phi}_{k}$ are i.i.d.
+sampled from $\mathcal{N}(0_{n\times 1},I_{n})$, and the measurement $y_{k}$ is given by $y_{k}=\phi_{k}\theta+v_{k}$, where $v_{k}\in{\mathbb{R}}^{p\times m}$ is the measurement noise and $\operatorname{vec}(v_{k})$ is i.i.d.
+sampled from $\mathcal{N}(0_{pm\times 1},\Sigma)$, where
+[…]
+and $\Sigma_{22}\triangleq 100\Sigma_{11}$.
+In this setup, the second column of the measurement noise has higher variance than that of the first column, and any two elements of the measurement noise are highly correlated with a correlation coefficient of $0.99$.
+We compare the performance of vec-permutation, column-by-column, and matrix update RLS.
+We set regularization terms as $\bar{\theta}_{0}=0_{nm\times 1}$ and $\bar{R}=I_{nm}$, $\theta_{j,0}=0_{n\times 1}$ and $R_{j}=I_{n}$ $j=1,2$, and $\theta_{0}=0_{n\times m}$ and $R=I_{n}$, respectively.
+For vec-permutation, we let $\bar{\Gamma}_{k}=\Sigma^{-1}$, giving the minimum variance estimator.
+For column-by-column, we let $\Gamma_{1,k}=\Sigma_{11}^{-1}$ and $\Gamma_{2,k}=\Sigma_{22}^{-1}$.
+Finally, for matrix update, we consider the three choices $\Gamma_{k}=I_{2}$, $\Gamma_{k}=\Sigma_{11}^{-1}$, and $\Gamma_{k}=\Sigma_{22}^{-1}$.
+[…]
+[…]
+6 Application to Online Identification for Indirect Adaptive Model Predictive Control
+Finally, a useful application of this work is efficient online identification for adaptive model predictive control. Consider a MIMO input-output system of the form
 $y_{k}=-\sum_{i=1}^{\hat{n}}F_{i}y_{k-i}+\sum_{i=0}^{\hat{n}}G_{i}u_{k-i},$ (41)
@@
 $\operatorname{vec}(\theta_{k+1})$ $=\operatorname{vec}(\theta_{k})+\bar{P}_{k+1}\bar{\phi}_{k}^{\rm T}(y_{k}-\bar{\phi}_{k}\operatorname{vec}(\theta_{k})).$ (51)
-Proof.
-This result follows from Lemma A.3. For further details, see equations (8) through (20) of [4]. ∎
+Proof 6.1.
+This result follows from Proposition 1. For further details, see equations (8) through (20) of [4].
 Next, we provide an alternate formulation using matrix RLS.
@@
 $\theta_{k+1}$ $=\theta_{k}+(y_{k}-\theta_{k}\phi_{k})\phi_{k}^{\rm T}P_{k+1}.$ (53)
-Proof.
-Note that, for all $k\geq 0$, $z_{k}(\hat{\theta})^{\rm T}z_{k}(\hat{\theta})=\operatorname{tr}(z_{k}(\hat{\theta})^{\rm T}z_{k}(\hat{\theta}))=\operatorname{tr}(z_{k}(\hat{\theta})z_{k}(\hat{\theta})^{\rm T})$.
-It then follows from (57) of Lemma A.2 that, for all $k\geq 0$, (43) can be written as
-$J_{k}(\hat{\theta})=\operatorname{tr}\Big{[}\sum_{i=0}^{k}z_{i}(\hat{\theta})z_{i}(\hat{\theta})^{\rm T}+(\hat{\theta}-\theta_{0}){P}_{0}^{-1}(\hat{\theta}-\theta_{0})^{\rm T}\Big{]}.$
-Proposition 5 then implies that, for all $k\geq 0$, (53) and $\theta_{k+1}^{\rm T}=\theta_{k}^{\rm T}+P_{k+1}\phi_{k}(y_{k}^{\rm T}-\phi_{k}^{\rm T}\theta_{k}^{\rm T})$ hold.
-Taking the transpose then yields (54).
-∎
-Example 1.
+Proof 6.2.
+This result follows from Proposition 3.
+Example 6.3.
 This example is from [10] and uses PCAC for the control of a flexible structure under harmonic and broadband disturbances.
@@
 Two actuators are placed at nodes 3 and 4 with control authority in the $x$-direction and $x$-direction displacement sensors are placed at nodes 5, 6, 7, and 8.
-The objective is to use PCAC to suppress the effects of harmonic and broadband disturbances without prior knowledge of the truss dynamics.
+The objective is to use PCAC to suppress the effects of broadband disturbances with constraints on actuator force and without prior knowledge of the truss dynamics.
 See [10] for further details.
@@
 Figure 3: Flexible truss structure from [10] with nodes labeled.
-TABLE III: Truss ex. computation time per step: mean $\pm$ std. deviation
+Table 3: Truss ex. computation time per step: mean $\pm$ std. deviation
 Vec-Permutation Matrix RLS Change
@@
 […]
-VI Conclusions
+7 Conclusions
 This work derives batch and recursive least squares algorithms for the identification of matrix parameters.
-Under the assumption of independent, identical column weighting, these methods minimize the same cost function as the vec-permutation approach.
+Under the assumption of independent, identical column weighting, this method minimizes the same cost function as the vec-permutation approach while significantly improving computational complexity.
 It is also shown how, under persistent excitation, convergence guarantees can be extended from the vector case to the matrix case.
 This approach can be used fast online identification of MIMO systems which is critical in indirect adaptive model predictive control.
+A future area of interest is studying how various RLS forgetting algorithms (e.g.
+[19]) can be applied to matrix update RLS.